For boolean algebra can you use an identity but make everything opposite? NOTs

Ex. x+!xy = x+y Can you do !x+x!y = !x+!y ???

or Ex. xy+!xy =y (minimization) Can you do !x!y + x!y = !y ???

Note this is boolean algebra. ! means NOT. Imagine a line above the particular letter to the right of it. \bar{x} <- like this, x bar.
Answer #1

Given r! X Y = x+y Exla =) Here Using distribution law in L.H.S (I + !x) (X+Y) - Ity Since (ht!x) = 1 [Unit Property] + 1(x+yGriven Ex-27 xx+10y=y (x+!x) = 1 by Unit Property (x +)y = y 1.Y=y I Ty=YT Similarly for !! Y 4 x ly = !Y 2+H;8. 1 X ! Y + x